# Diff-Reg v1: Diffusion Matching Model for Registration Problem

### v1

@@ Abstract
Establishing reliable correspondences is essential for registration tasks such as 3D and 2D3D registration. Existing methods commonly leverage geometric or semantic point features to generate potential correspondences. However, these features may face challenges such as large deformation, scale inconsistency, and ambiguous matching problems (e.g., symmetry). Additionally, many previous methods, which rely on single-pass prediction, may struggle with local minima in complex scenarios. To mitigate these challenges, we introduce a diffusion matching model for robust correspondence construction. Our approach treats correspondence estimation as a denoising diffusion process within the doubly stochastic matrix space, which gradually denoises (refines) a doubly stochastic matching matrix to the ground-truth one for high-quality correspondence estimation. It involves a forward diffusion process that gradually introduces Gaussian noise into the ground truth matching matrix and a reverse denoising process that iteratively refines the noisy matching matrix. In particular, the feature extraction from the backbone occurs only once during the inference phase. Our lightweight denoising module utilizes the same feature at each reverse sampling step. Evaluation of our method on both 3D and 2D3D registration tasks confirms its effectiveness. The code will be made available online.
@@ III-D Diffusion Model in Doubly Stochastic Matrix space
$\tilde{E}^{t}=\text{sigmoid}(E^{t}),\quad\tilde{E}^{t}=\mathbf{f}_{\text{sinkhorn}}(E^{t})$ (2)
where the $\mathbf{f}_{\text{sinkhorn}}$ operation is from the Sinkhorn algorithm[50].
@@ III-E The Lightweight Denoising Module $g_{\theta}$
Algorithm 1 Denoising Module $g_{\theta}$ in 3D Registration Task.
1:Sampled matching matrix $E^{t}\in\mathbb{R}^{N\times M}$; Point clouds $P,Q\in\mathbb{R}^{3}$ and associated point features $F^{\hat{P}},F^{\hat{Q}}$.
2:Target matching matrix $\hat{E}_{0}$.
3:function $g_{\theta}$($E^{t},\hat{P},\hat{Q},F^{\hat{P}},F^{\hat{Q}}$)
4:     $\tilde{E}_{t}\leftarrow\mathbf{f_{sinkhorn}}(E^{t})$
5:     $\hat{R}_{t},\hat{t}_{t}\leftarrow\mathbf{soft\_procrustes}(\tilde{E}_{t},\hat{P},\hat{Q})$
6:     $\hat{P}_{t}\leftarrow\mathbf{warping}(P,\hat{R}_{t},\hat{t}_{t})$
7:     $\tilde{F}^{\hat{P}_{t}},\tilde{F}^{\hat{Q}_{t}}\leftarrow f_{\theta}(\hat{P}_{t},\hat{Q},F^{\hat{P}},F^{\hat{Q}},\Theta(\hat{P}_{t}),\Theta(\hat{Q}))$
8:     $\tilde{E}_{0}\leftarrow\mathbf{matching\_logits}(\tilde{F}^{\hat{P}_{t}},\tilde{F}^{\hat{Q}_{t}},\Theta(\hat{P}_{t}),\Theta(\hat{Q}))$
9:     $\hat{E}_{0}\leftarrow\mathbf{f_{sinkhorn}}(\tilde{E}_{0})$
10:     return $\hat{E}_{0}$
11:end function
@@ IV Experiments
1:Point clouds $\hat{P}$, $\hat{Q}\in\mathbb{R}^{3}$ and associated point features $F^{\hat{P}},F^{\hat{Q}}$.
@@
3:     Sample $E^{0}\sim q(E^{0})$
4:     $N\times M\leftarrow E^{0}.\text{shape}$
@@
6:     $\epsilon\sim N(0,1)^{N\times M}$
7:     $\tilde{E}^{t}\leftarrow\sigma(\sqrt{\bar{\alpha}_{t}}E^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon)$
8:     $\hat{E}_{0}\leftarrow g_{\theta}(\tilde{E}^{t},\hat{P},\hat{Q},F^{\hat{P}},F^{\hat{Q}})$
9:     Optimize $L_{t}=\text{Focal\_loss}(\hat{E}_{0},E^{0})$
10:end while
@@
1:Initial matching matrix $E^{T}$ from backbone or white noise; Point clouds $\hat{P},\hat{Q}\in\mathbb{R}^{3}$ and associated point features $F^{\hat{P}},F^{\hat{Q}}$.
2:Target matching matrix $E^{0}$.
3:$N\times M\leftarrow E^{T}.\text{shape}$
@@
5:     ${z_{t}}\sim N(0,1)^{N\times M}$ if $t>1$ else ${z_{t}}\leftarrow 0^{N\times M}$
6:     $\hat{E}_{0}\leftarrow g_{\theta}(\hat{E}^{t},\hat{P},\hat{Q},F^{\hat{P}},F^{\hat{Q}})$
7:     $\epsilon_{t}\leftarrow\frac{\hat{E}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{E}^{t}$
9:     $\hat{E}^{t-1}\leftarrow\sqrt{\alpha_{t-1}}\hat{E}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{t}+\sigma_{t}z_{t}$
10:end for
@@ Results.
Algorithm 4 Training Diff-Reg in 2D3D Registration Task
1:Coarse level image and points $\hat{X}\in\mathbb{R}^{\hat{H}{\times}{\hat{W}}{\times}{3}}$, $\hat{Q}\in\mathbb{R}^{3}$ and associated features $F^{\hat{X}},F^{\hat{X}}_{dino},F^{\hat{Q}}$.
@@
3:     Sample $E^{0}\sim q(E^{0})$
4:     $N\times M\leftarrow E^{0}.\text{shape}$
@@
6:     $\epsilon\sim N(0,1)^{N\times M}$
7:     $\tilde{E}^{t}\leftarrow\sigma(\sqrt{\bar{\alpha}_{t}}E^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon)$
8:     $\hat{E}_{0}\leftarrow g_{\theta}(\tilde{E}^{t},\hat{X},\hat{Q},F^{\hat{X}},F^{\hat{X}}_{dino},F^{\hat{Q}})$
9:     Optimize $L_{t}=\text{Focal\_loss}(\hat{E}_{0},E^{0})$
@@
Algorithm 5 Sampling by Diff-Reg in 2D3D Registration Task
1:Initial matching matrix $E^{T}$ from backbone or white noise; Coarse level image and points $\hat{X}\in\mathbb{R}^{\hat{H}{\times}{\hat{W}}{\times}{3}}$, $\hat{Q}\in\mathbb{R}^{3}$ and associated features $F^{\hat{X}},F^{\hat{X}}_{dino},F^{\hat{Q}}$.
2:Target matching matrix $E^{0}$.
3:$N\times M\leftarrow E^{T}.\text{shape}$
@@
5:     ${z_{t}}\sim N(0,1)^{N\times M}$ if $t>1$ else ${z_{t}}\leftarrow 0^{N\times M}$
6:     $\hat{E}_{0}\leftarrow g_{\theta}(\tilde{E}^{t},\hat{X},\hat{Q},F^{\hat{X}},F^{\hat{X}}_{dino},F^{\hat{Q}})$
7:     $\epsilon_{t}\leftarrow\frac{\hat{E}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{E}^{t}$
@@
9:     $\hat{E}^{t-1}\leftarrow\sqrt{\alpha_{t-1}}\hat{E}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{t}+\sigma_{t}z_{t}$

### v2

@@ Abstract
Establishing reliable correspondences is essential for registration tasks such as 3D and 2D3D registration. Existing methods commonly leverage geometric or semantic point features to generate potential correspondences. However, these features may face challenges such as large deformation, scale inconsistency, and ambiguous matching problems (e.g., symmetry). Additionally, many previous methods, which rely on single-pass prediction, may struggle with local minima in complex scenarios. To mitigate these challenges, we introduce a diffusion matching model for robust correspondence construction. Our approach treats correspondence estimation as a denoising diffusion process within the doubly stochastic matrix space, which gradually denoises (refines) a doubly stochastic matching matrix to the ground-truth one for high-quality correspondence estimation. It involves a forward diffusion process that gradually introduces Gaussian noise into the ground truth matching matrix and a reverse denoising process that iteratively refines the noisy matching matrix. In particular, the feature extraction from the backbone occurs only once during the inference phase. Our lightweight denoising module utilizes the same feature at each reverse sampling step. Evaluation of our method on both 3D and 2D3D registration tasks confirms its effectiveness. The code is available at https://github.com/wuqianliang/Diff-Reg.
@@ III-D Diffusion Model in Doubly Stochastic Matrix space
[…]
where the $\mathbf{f}_{\text{sinkhorn}}$ operation is from the Sinkhorn algorithm [50] and $f_{\epsilon}=(\epsilon\%1)(abs(\epsilon)/\epsilon)\eta$. We empirically set $\eta=1.5$ and $\epsilon_{0}\sim\mathcal{N}(\epsilon;0,\textbf{I})$.
@@ III-E The Lightweight Denoising Module $g_{\theta}$
Algorithm 1 Denoising Module $g_{\theta}$ for 3D Registration Task.
1:Sampled matching matrix $\mathbf{E}^{t}\in\mathbb{R}^{N\times M}$; Point clouds $\mathbf{P},\mathbf{Q}\in\mathbb{R}^{3}$ and associated point features $\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:Target matching matrix $\hat{\mathbf{E}}_{0}$.
3:function $g_{\theta}$($\mathbf{E}^{t},\hat{\mathbf{P}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}}$)
4:     $\tilde{\mathbf{E}}_{t}\leftarrow\mathbf{f_{sinkhorn}}(\mathbf{E}^{t})$
5:     $\mathbf{\hat{R}}_{t},\mathbf{\hat{t}}_{t}\leftarrow\mathbf{soft\_procrustes}(\tilde{\mathbf{E}}_{t},\hat{\mathbf{P}},\hat{\mathbf{Q}}),\ \hat{\mathbf{P}}_{t}\leftarrow\mathbf{warping}(\mathbf{P},\mathbf{\hat{R}}_{t},\mathbf{\hat{t}}_{t})$
6:     $\tilde{\mathbf{F}}^{\hat{\mathbf{P}}_{t}},\tilde{\mathbf{F}}^{\hat{\mathbf{Q}}_{t}}\leftarrow f_{\theta}(\hat{\mathbf{P}}_{t},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}},\Theta(\hat{\mathbf{P}}_{t}),\Theta(\hat{\mathbf{Q}}))$
7:     $\tilde{\mathbf{E}}_{0}\leftarrow\mathbf{matching\_logits}(\tilde{\mathbf{F}}^{\hat{\mathbf{P}}_{t}},\tilde{\mathbf{F}}^{\hat{\mathbf{Q}}_{t}},\Theta(\hat{\mathbf{P}}_{t}),\Theta(\hat{\mathbf{Q}}))$
8:     $\hat{\mathbf{E}}_{0}\leftarrow\mathbf{f_{sinkhorn}}(\tilde{\mathbf{E}}_{0})$
9:     return $\hat{\mathbf{E}}_{0}$
10:end function
@@ IV Experiments
1:Point clouds $\hat{\mathbf{P}}$, $\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated point features $\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}}$.
@@
3:     Sample $\mathbf{E}^{0}\sim q(\mathbf{E}^{0})$
4:     $N\times M\leftarrow\mathbf{E}^{0}.\text{shape}$
@@
6:     $\epsilon\sim\mathcal{N}(0,1)^{N\times M}$
7:     if Rigid then
8:         $\mathbf{E}^{t}\leftarrow\sqrt{\bar{\alpha}_{t}}\mathbf{E}^{0}+\sqrt{1-\bar{\alpha}_{t}}f_{\epsilon}(\epsilon),\quad\tilde{\mathbf{E}}^{t}=\mathbf{E}^{t}-\rm{Min}(\mathbf{E}^{t})$
9:     else
10:         $\mathbf{E}^{t}=\sqrt{\bar{\alpha_{t}}}\mathbf{E}^{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon_{0},\quad\tilde{\mathbf{E}}^{t}=\rm{Sigmoid}(\mathbf{E}^{t})$
11:     end if
12:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\tilde{\mathbf{E}}^{t},\hat{\mathbf{P}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}})$
13:     Optimize $L_{t}=\rm{Focal\_loss}(\hat{\mathbf{E}}_{0},\mathbf{E}^{0})$
14:end while
@@
1:Initial matching matrix $E^{T}$ from backbone or white noise; Point clouds $\hat{\mathbf{P}},\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated point features $\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:Target matching matrix $\tilde{\mathbf{E}}^{0}$.
3:$N\times M\leftarrow\mathbf{E}^{T}.\text{shape}$
@@
5:     ${\mathbf{z}_{t}}\sim N(0,1)^{N\times M}$ if $t>1$ else ${\mathbf{z}_{t}}\leftarrow\mathbf{0}^{N\times M}$
6:     if Rigid then
7:         $\tilde{\mathbf{E}}^{t}=\mathbf{E}^{t}-\rm{Min}(\mathbf{E}^{t})$
8:     end if
9:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\hat{\mathbf{E}}^{t},\hat{\mathbf{P}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}})$
10:     $\epsilon_{t}\leftarrow\frac{\hat{\mathbf{E}}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\mathbf{E}}^{t}$
11:     $\sigma_{t}\leftarrow\sqrt{\frac{(1-\alpha_{t-1})}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}$
12:     $\hat{\mathbf{E}}^{t-1}\leftarrow\sqrt{\alpha_{t-1}}\hat{\mathbf{E}}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{t}+\sigma_{t}\mathbf{z}_{t}$
13:end for
14:if Rigid then
15:     $\tilde{\mathbf{E}}^{0}=\tilde{\mathbf{E}}^{0}-\rm{Min}(\tilde{\mathbf{E}}^{0}),\quad\tilde{\mathbf{E}}^{0}=\mathbf{f}_{\rm{sinkhorn}}(\tilde{\mathbf{E}}^{0})$
16:else
17:     $\tilde{\mathbf{E}}^{0}=\rm{Sigmoid}(\tilde{\mathbf{E}}^{0})$
18:end if
@@ Results.
Algorithm 4 Training Diff-Reg in 2D-3D Registration Task
1:Coarse level image and points $\hat{\mathbf{X}}\in\mathbb{R}^{\hat{H}{\times}{\hat{W}}{\times}{3}}$, $\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated features $\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}}$.
@@
3:     Sample $\mathbf{E}^{0}\sim q(\mathbf{E}^{0})$
4:     $N\times M\leftarrow\mathbf{E}^{0}.\text{shape}$
@@
6:     $\epsilon\sim\mathcal{N}(0,1)^{N\times M}$
7:     $\mathbf{E}^{t}\leftarrow\sqrt{\bar{\alpha}_{t}}\mathbf{E}^{0}+\sqrt{1-\bar{\alpha}_{t}}f_{\epsilon}(\epsilon),\quad\tilde{\mathbf{E}}^{t}=\mathbf{E}^{t}-\rm{Min}(\mathbf{E}^{t})$
8:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\tilde{\mathbf{E}}^{t},\hat{\mathbf{X}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}})$
9:     Optimize $L_{t}=\text{Focal\_loss}(\hat{\mathbf{E}}_{0},\mathbf{E}^{0})$
@@
Algorithm 5 Sampling by Diff-Reg in 2D-3D Registration Task
1:Initial matching matrix $\mathbf{E}^{T}$ from backbone or white noise; Coarse level image and points $\hat{\mathbf{X}}\in\mathbb{R}^{\hat{H}{\times}{\hat{W}}{\times}{3}}$, $\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated features $\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:Target matching matrix $\tilde{\mathbf{E}}^{0})$.
3:$N\times M\leftarrow\mathbf{E}^{T}.\text{shape}$
@@
5:     ${\mathbf{z}_{t}}\sim\mathcal{N}(0,1)^{N\times M}$ if $t>1$ else ${\mathbf{z}_{t}}\leftarrow\mathbf{0}^{N\times M}$
6:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\tilde{\mathbf{E}}^{t},\hat{\mathbf{X}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}})$
7:     $\epsilon_{t}\leftarrow\frac{\hat{\mathbf{E}}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\mathbf{E}}^{t}$
@@
9:     $\hat{\mathbf{E}}^{t-1}\leftarrow\sqrt{\alpha_{t-1}}\hat{\mathbf{E}}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{t}+\sigma_{t}\mathbf{z}_{t}$
@@
11:$\tilde{\mathbf{E}}^{0}=\mathbf{f}_{\rm{sinkhorn}}(\tilde{\mathbf{E}}^{0})$